### v1

# Invariants of Superalgebras as Complex Integrals

Abstract
In [1] we defined integrals that approximated the Poincaré series of the invariants and cocomiants of the general linear Lie supergroup or superalgebra. Budzik suggested in [3] a way to adapt this method to get the exact Poincaré series. The purpose of this paper is to prove that Budzik’s ideas are correct. As a consequence we prove that the Poincaré series are rational functions.
1 Budzik’s Conjecture
The basic objects of that form the subject of this paper can be described in several different ways. The proof of our main theorem will mosty focus on their definition based on the character theory of the symmetic group, but in order to provide motivation and context, we also mention some of the connections to trace identities and to invariant theory.
We will use greek letters $\lambda,$ $\mu$ and $\nu$ to denote partitions of some unspecified, common integer $n$. When we specify the parts of a partition $\lambda=(\lambda_{1},\lambda_{2},\ldots)$, we will always intend the parts to be in non-increasing order $\lambda_{1}\geq\lambda_{2}\geq\cdots$. For each partition $\lambda$ of $n$, we let $\chi^{\lambda}$ be the corresponding character of the symmetric group $S_{n}$. Given $\chi^{\mu}$ and $\chi^{\nu}$ the inner tensor product $\chi^{\mu}\otimes\chi^{\nu}$ can be decomposed into a sum of irreducible parts, whose multiplicities we denoted by $\gamma_{\mu,\nu}^{\lambda}$, so that $\chi^{\mu}\otimes\chi^{\nu}=\sum\gamma_{\mu,\nu}^{\lambda}\chi^{\lambda}$.
Let $k$ and $\ell$ be two fixed non-negative integers, at least one of which is positive. Then $H(k,\ell),$ the $k\times\ell$ hook, is defined to be the set of partitions $\lambda$ with at most $k$ parts greater than or equal to $\lambda$, i.e., such that $\lambda_{k+1}\leq\ell$. The basic objects we will be studying are the multiplicities occuring are the tensor products $\sum_{\mu\in H(k,\ell)}\chi^{\mu}\otimes\chi^{\mu}$. The multiplicity $m_{\lambda}$ of $\chi^{\lambda}$ in such a sum is given by
$m_{\lambda}=\sum_{\mu\in H(k,\ell)}\gamma_{\mu,\mu}^{\lambda}.$ (1)
Sometimes we will write $m_{\lambda}(k,\ell)$ if we need to specify the $k$ and $\ell$. We associate to $m_{\lambda}$ the Poincaré series $P(k,\ell;n)=\sum m_{\lambda}s_{\lambda}(t_{1},\ldots,t_{n})$, where $s_{\lambda}$ is the Schur function, and more generally,
$P(k,\ell;n,m)=\sum m_{\lambda}HS_{\lambda}(t_{1},\ldots,t_{n};u_{1},\ldots,u_{m}),$
where $HS_{\lambda}$ is the hook Schur function, defined below for the reader not familiar with them.
The connection to trace identities involves the Kemer algebras. Let $E$ be an infinite dimensional Grassmann algebra in characteristic zero. This algebra has a natural $\mathbb{Z}_{2}$-grading, $E=E_{0}\oplus E_{1}$ with respect to which the degree 0 elements are central and the degree 1 elements anticommute with each other. The Kemer algebra, $M_{k,\ell}$, is defined to be the set of $(k+\ell)\times(k+\ell)$ matrices with entries from $E$ of the form $\left(\begin{smallmatrix}A&B\\ C&D\end{smallmatrix}\right)$ such that $A$ is a $k\times k$ matrix with entries from $E_{0}$, $D$ is an $\ell\times\ell$ matrix with entries from $E_{0}$, and $B$ and $C$ have entries from $E_{1}$. This algebra has a natural trace function $tr:M_{k,\ell}\rightarrow E_{0}$ which takes a matrix as above to tr(A)tr(D)tr(A)-tr(D).
To construct generic element for $M_{k,\ell}$ we use the free supercommutative algebra on the degree 0 variables $x_{ij}^{(\alpha)}$ and the degree 1 variables $e_{ij}^{(\alpha)}$. Then $X_{\alpha}$ is defined to be the matrix $\left(\begin{smallmatrix}A&B\\ C&D\end{smallmatrix}\right)$ in which the dimensions of $A,B,C,D$ are as above, and the $(i,j)$ entries of $A$ and $D$ are $x_{ij}^{(\alpha)},$ and the $(i,j)$ entries of $B$ and $D$ are $e_{ij}^{(\alpha)}$. One next constructs $F[X_{1},\ldots,X_{n}]$, the algebra generated by $n$ such matrices and $TR(k,\ell;n)$, the algebra generated by the traces of the elements of $F[X_{1},\ldots,X_{n}]$. The algebra $TR(k,\ell;n)$ has a natural $n$-fold grading and so a Poincaré series in $n$ variables. It turns out that this Poincaré series is none other than $P(k,\ell;n)$.
This Poincaré series is also the Poincaré series for the invariants of the general linear Lie superalgebra or supergroup. In deference to [3] we focus on the supergroup, $U(k|\ell)$ (also denoted $PL(k,\ell)$), the group of invertible matrices in $M_{k,\ell}$. This group acts on $M_{k,\ell}$ by conjugation and, by extension, on the ring of $E_{0}$-linear maps $M_{k,\ell}^{n}\rightarrow E$. Such a function $F$ is said to be invariant if
$F(ga_{1}g^{-1},\ldots,ga_{n}g^{-1})=F(a_{1},\ldots,a_{n}),$
for all $a_{1},\ldots,a_{n}\in M_{k,\ell}$ and all $g\in U(k|\ell)$. The algebra of invariant functions on $M_{k,\ell}^{n},$ polynomial in the entries, also has an $n$-fold grading and also has Poincaré series $P(k,\ell,n)$. In fact it is generated by the trace polynomials.
The series $P(k,\ell;n,m)$ also has a natural interpretation. We refer the interested reader to [1]. We also mention that we learned from [3] that these series have signifigance in physics: “These arise in physics e.g. as free partition functions of $U(N|M)$ gauge theories or supersymmetric indices.”
In [1] we studied these Poincaré series using the complex integrals. Let $\Delta$ (not the same as the $\Delta$ from [1]) be defined as
$\Delta=\frac{\prod_{1\leq i\neq j\leq k}(1-x_{i}x_{j}^{-1})\prod_{1\leq i\neq j\leq\ell}(1-y_{i}y_{j}^{-1})}{\prod_{i=1}^{k}\prod_{j=1}^{\ell}(1+x_{i}y_{j}^{-1})(1+x_{i}^{-1}y_{j})}.$
Let $X=\{x_{1},\ldots,x_{k}\},\ Y=\{y_{1},\ldots,y_{k}\}$ and let $Z_{0}$ be the set of variables $XX^{-1}\cup YY^{-1}$ and let $Z_{1}$ be the set of variables $XY^{-1}\cup X^{-1}Y$. Then for each $\lambda$ we define $m_{\lambda}^{\prime}$ or $m_{\lambda}^{\prime}(k,\ell)$ via
$m_{\lambda}^{\prime}=\frac{1}{k!\ell!(2\pi i)^{k+\ell}}\oint_{|y_{i}|=1}\oint_{|x_{i}|=1+\epsilon}HS_{\lambda}(Z_{0};Z_{1})\Delta\frac{dx_{1}}{x_{1}}\wedge\cdots\wedge\frac{dy_{\ell}}{y_{\ell}},$ (2)
where $HS_{\lambda}$ is the hook Schur function, and the $\epsilon>0$ is to make sure that all of the poles at $x_{i}=y_{j}$ lie inside of the inner integrals. In [1] we proved that $m_{\lambda}=m_{\lambda}^{\prime}$ for all $\lambda$ occuring with non-zero multiplicity in $\sum_{\mu\in H(k,\ell)}\chi^{\mu}\otimes\chi^{\mu}$, but not in $\sum_{\mu\in H(k^{\prime},\ell,)}\chi^{\mu}\otimes\chi^{\mu},$ for any smaller hook $H(k^{\prime},\ell^{\prime})$. Based on this, Budzik conjectured in [3] (slightly reformulated here) that
Budzik’s Conjecture.
$m_{\lambda}^{\prime}(k,\ell)=m_{\lambda}(k,\ell)-m_{\lambda}(k-1,\ell-1).$
By induction, this would imply that
$m_{\lambda}(k,\ell)=\sum_{i=0}^{\min(k,\ell)}m_{\lambda}^{\prime}(k-i,\ell-i).$ (3)
Likewise, setting $P^{\prime}(k,\ell;n)$ equal to $\sum m_{\lambda}^{\prime}S_{\lambda}(t_{1},\ldots,t_{n})$, Budzik’s conjecture would imply that $P^{\prime}(k,\ell;n)=P(k,\ell;n)-P(k-1,\ell-1;n)$ and so
$P(k,\ell;n)=\sum_{i=0}^{\min(k,\ell)}P^{\prime}(k-i,\ell-i;n)$ (4)
More generally, setting $P^{\prime}(k,\ell;n,m)$ equal to $\sum m_{\lambda}^{\prime}HS(t_{1},\ldots,t_{n};u_{1},\ldots,u_{m})$ we would have $P^{\prime}(k,\ell;n,m)=P(k,\ell;n,m)-P(k-1,\ell-1k;n)$ and
$P(k,\ell;n,m)=\sum_{i=0}^{\min(k,\ell)}P^{\prime}(k-i,\ell-i;n,m).$ (5)
Our main purpose in this paper is to prove Budzik’s conjecture and, hence, all of the above equations.
Note that (2) together with Budzik’s conjection in the form of (3) allows the computation of any $m_{\lambda}$. Likewise, $P(k,\ell;n)$ and $P(k,\ell;n,m)$ can be computed using equations (4) and (5) using Theorem 2.4 from [1]:
Theorem 1.
$P^{\prime}(k,\ell;n,m)$ equals $\frac{1}{k!\ell!(2\pi i)^{k+\ell}}$ times the integral
$\oint_{|y_{i}|=1}\oint_{|x_{i}|=1+\epsilon}\frac{\prod_{z\in Z_{0}}\prod_{i}(1+zu_{i})\prod_{i}\prod_{z\in Z_{1}}(1+zt_{i})}{\prod_{i}\prod_{z\in Z_{0}}(1-zt_{i})\prod_{i}\prod_{z\in Z_{1},}(1-zu_{i})}\Delta\frac{dx_{1}}{x_{1}}\wedge\cdots\wedge\frac{dy_{\ell}}{y_{\ell}}$
and, taking $Y=\emptyset$, $P^{\prime}(k,\ell;n)$ equals $\frac{1}{k!\ell!(2\pi i)^{k+\ell}}$ times the integral
$\oint_{|y_{i}|=1}\oint_{|x_{i}|=1+\epsilon}\frac{\prod_{i}\prod_{z\in Z_{1}}(1+zt_{i})}{\prod_{i}\prod_{z\in Z_{0}}(1-zt_{i})\prod_{i}}\Delta\frac{dx_{1}}{x_{1}}\wedge\cdots\wedge\frac{dy_{\ell}}{y_{\ell}}$
Before leaving this section we also mention that Budzik computed $P^{\prime}(k,\ell;1)$ for all $k$ and $\ell$, namely,
$P^{\prime}(k,\ell;1)=\frac{t^{k\ell}}{\prod_{i=1}^{k}(1-t^{i})\prod_{j=1}^{\ell}(1-t^{j})}.$
2 Properties of Hook Schur Functions
We fix two sets of variables $X=\{x_{1},\ldots,x_{k}\}$ and $Y=\{y_{1},\ldots,y_{k}\}$. For each partition $\lambda$ we let $HS_{\lambda}(X;Y)$ be the hook Schur function on $\lambda$. This is commonly defined as $\sum_{\mu\subseteq\lambda}s_{\lambda}(X)s_{\lambda/\mu}(Y)$, where $s$ is the ordinary Schur function. Some authors study the super Schur funcitons $s_{\lambda}(X/Y)$ instead of the hook Schur functions $HS_{\lambda}(X;Y)$. Each has its advantages, but they are essentially equivalent, since $HS_{\lambda}(X;-Y)=s_{\lambda}(X/Y)$.
We will need these properties of hook Schur funtions:
Theorem 2 (The Hook Theorem).
$HS_{\lambda}(X;Y)=0$ unless $\lambda\in H(k,\ell).$
Theorem 3.26 of [2] implies that if $\mu,\nu\in H(k,\ell)$ then $\gamma_{\mu,\nu}^{\lambda}=0$, and so $m_{\lambda}=0$ unless $\lambda\in H(k^{2}+\ell^{2},2k\ell)$. This means that $P(k,\ell;k^{2}+\ell^{2};2k\ell)$ encodes complete information about the $m_{\lambda}$, and no $P(k,\ell;n,m)$ does for smaller $n,m$.
A partition in $H(k,\ell)$ with $\lambda_{k}\geq\ell$ will not belong to $H(k-1,\ell-1)$. Such a partition is called typical and the set of such is denoted $H^{\prime}(k,\ell)$.
The second theorem is due to Józefiak and Pragacz, [4]. It can also be found in [5], 1.3 Ex. 24. To each partition $\lambda$ we say $(i,j)\in\lambda$ if $\lambda_{i}\leq j$, or, more geometrically, if the Young tableau associated to $\lambda$ has a box in row $i$, column $j$. Then $f_{\lambda}$ is defined to be the product $\prod(x_{i}+y_{j})$ over all $(i,j)\in\lambda$, with the convention that $x_{i}=0$ and $y_{j}=0$ for $i>k$ or $j>\ell$. For example, if $(k,\ell)=(2,1)$ and $\lambda=(2,1)$, then $f_{\lambda}=(x_{1}+y_{1})x_{1}^{2}(x_{2}+y)$. Józefiak and Pragacz’s theorem gives an alternate description of hook Schur functions.
Theorem 3 (Józefiak and Pragacz).
$HS_{\lambda}(X;Y)=$
$\sum_{\sigma\in S_{k}\times S_{\ell}}\sigma\left[\frac{f_{\lambda}(X;Y)}{\prod_{i<j}(1-x_{j}x_{i}^{-1})\prod_{i<j}(1-y_{j}y_{i}^{-1})}\right]$
So, for the example given above $HS_{(2,1)}(x_{1},x_{2};y_{1})=$
$\frac{(x_{1}+y_{1})x_{1}(x_{2}+y_{1})}{1-x_{2}x_{1}^{-1}}+\frac{(x_{2}+y_{1})x_{2}(x_{1}+y_{1})}{1-x_{1}x_{2}^{-1}}=(x_{1}+y_{1})(x_{2}+y_{1})(x_{1}+x_{2}).$
A weaker version of Józefiak and Pragacz’s theorem is called the factorization theorem, and it is due to Berele and Regev from [2].
Theorem 4 (Berele and Regev).
Let $\lambda$ be a typical partition, and define the partitions $\alpha$ and $\beta$, via $\alpha=(\lambda_{1}-\ell,\ldots,\lambda_{k}-\ell)$ and $\beta=(\lambda_{1}^{\prime}-k,\ldots,\lambda_{\ell}^{\prime}-k)$. Then $HS_{\lambda}(X;Y)=\prod_{i,j}(x_{i}+y_{j})s_{\alpha}(X)s_{\beta}(Y)$.
The next theorem is due to Rosas in [7]. It gives a hook Schur function interpretation of the coefficients $\gamma_{\mu\nu}^{\lambda}$ occuring in the tensor product of the symmetric group characters.
Theorem 5 (Rosas).
Given 4 sets of variables $X$, $Y$, $T$ and $U$,
$HS_{\lambda}(XT,YU;XU,YT)=\sum_{\mu,\nu}\gamma_{\mu,\nu}^{\lambda}HS_{\mu}(X;Y)HS_{\nu}(T;U).$
In the case of $U=\emptyset$ this says that
$HS_{\lambda}(XT;YT)=\sum_{\mu,\nu}\gamma_{\mu,\nu}^{\lambda}HS_{\mu}(X;Y)s_{\nu}(T).$
Finally, we will need this hook generalization of Cauchy’s identity found in [6]
Theorem 6 (Berele and Remmel).
Given 4 sets of variables $X$, $Y$, $T$ and $U$,
$\sum_{\lambda}HS_{\lambda}(X;Y)HS_{\lambda}(T;U)=\frac{\prod(1+x_{i}u_{j})\prod(1+y_{i}t_{j})}{\prod(1-x_{i}t_{j})\prod(1-y_{i}u_{j})},$
where the $x_{i}$ run over the elements of $X$, etc.
3 The Inner Product
There is an inner product on symmetric functions in $n$ variables with respect to which the Schur functions form an orthonormal basis. It can be defined using complex integrals via $\langle f,g\rangle=(2\pi i)^{-n}(n!)^{-1}$ times
$\oint_{|z_{1}|=1}\cdots\oint_{|z_{n}|=1}f(z_{1},\ldots,z_{n})g(z_{1}^{-1},\ldots,z_{n}^{-1})\prod_{i\neq j}(1-z_{i}z_{j}^{-1})\frac{dz_{1}}{z_{1}}\wedge\cdots\wedge\frac{dz_{n}}{z_{n}}.$
We now define a new inner product, related to but not the same as the inner product from [1]. In this setting, $f$ and $g$ will be functions of two sets of variables $X$ and $Y$, with $|X|=k$ and $|Y|=\ell$. This new inner product will be defined via $\langle f,g\rangle=(2\pi i)^{-(k+\ell)}(k!)^{-1}(\ell!)^{-1}$ times the integral
$\oint_{|y|=1}\oint_{|x|=1+\epsilon}f(X;Y)g(X^{-1};Y^{-1})\Delta\frac{dx_{1}}{x_{1}}\wedge\cdots\wedge\frac{dy_{\ell}}{y_{\ell}}.$
Under this product, the hook Schur functions are not orthonormal. We proved in [1] as an easy consequence of the factorization theorem, that if $\lambda$ and $\mu$ are typical then $\langle HS_{\mu},HS_{\nu}\rangle=\delta_{\mu,\nu}$. The main result of this section is the following.
Lemma 7.
If $\mu$ and $\nu$ are not both in $H^{\prime}(k,\ell)$, then $\langle HS_{\mu},HS_{\nu}\rangle=0$.
If $\mu$ or $\nu$ are not in $H(k,\ell)$, then $H_{\mu}(X;Y)$ or $H_{\nu}(X;Y)$ will be zero, so we will focus on $\mu,\nu\in H(k,\ell)$. In order to prove this lemma, we need the folowing general fact about integrals of rational functions. Just as for polynomials and rational functions, the degree of a Laurent polynomial is defined to be the degree of the highest degree term, and the degree of a quotient of Laurent polynomials is the difference between the degrees of the numerator and denominator.
Lemma 8.
Let $f=f(z_{1},\ldots,z_{n})$ and $g=g(z_{1},\ldots,z_{n})$ be Laurant polynomials with leading terms of degrees $a$ and $b$, respectively, with a<b1a<b-1. Let $|z|=r$ be a circle containing all of the roots of $g(z)$. Then $\oint_{|z|=r}f(z)/g(z)dz=0$.
Proof.
By Cauchy’s residue theorem $\oint_{|z|=r}f(z)/g(z)dz=0$ is not changed if we replace the circle $|z|=r$ with a larger circle $|z|=R$. For large $R$ when $|z|=R$, $|f(z)/g(z)|$ is $O(R^{a-b})$. Since the circumfrence of this circle is $2\pi R$, the absolute value of the integral is $O(R^{a-b+1})$, which tends to zero. ∎
We now analyze the fraction
$\frac{HS_{\mu}(X;Y)HS_{\nu}(X^{-1};Y^{-1})\prod(1-x_{i}x_{j}^{-1})\prod(1-y_{i}y_{j}^{-1})}{\prod(1-x_{i}y_{j}^{-1})(1-x_{i}^{-1}y_{j})\prod x_{i}\prod y_{j}}$ (6)
Lemma 9.
If $\mu$ is not typical, then the integrand (6) can be written as a sum of terms of the form $x^{\alpha}y^{\beta}\prod(x_{i}+y_{j})^{-2}$, each of degree at most 2-2 in one of the $x_{i}$.
Proof.
Isolating the term
$\frac{\prod(1-x_{i}x_{j}^{-1})\prod(1-y_{i}y_{j}^{-1})}{\prod(1-x_{i}y_{j}^{-1})(1-x_{i}^{-1}y_{j})\prod x_{i}\prod y_{j}},$ (7)
the numerator is of degree k1k-1 in each $x_{i}$, and the denominator is of degree $\ell+1$, so the degree of the fraction is $k-\ell-2$ in each $x_{i}$.
By Józefiak and Pragacz’s Theorem, $HS_{\mu}(X;Y)$ is a sum of the terms
$\sigma\left[f_{\mu}(X,Y)\prod_{i<j}(1-x_{j}x_{i}^{-1})^{-1}\prod_{i<j}(1-y_{j}y_{i}^{-1})^{-1}\right].$ (8)
Since $\mu$ is not typical, $\mu_{k}\leq\ell-1$, so $f_{\mu}(X;Y)$ is of degree at most $\ell-1$ in $x_{k}$; also $\prod_{i<j}(1-x_{j}x_{i}^{-1})^{-1}$ will be of degree k1k-1 in $x_{k}$, so (8) will be of degree at most $\ell-k$ in $x_{k}$, and (6) will be of degree at most 2. Under a permutation $\sigma$, $f_{\lambda}(X;Y)/\prod(1-x_{j}x_{i}^{-1})$ will be of degree at most $\ell-k$ in $x_{\sigma(k)}$, and the product with (7) at most 2-2. ∎
Proof of Lemma 7.
By the previous lemma, if $\mu$ is not typical, then $\langle HS_{\mu}(X;Y)HS_{\nu}(X;Y)\rangle$ is a linear combination of integrals of the form
$\oint_{|y|=1}\oint_{|x|=1+\epsilon}\frac{x^{\alpha}y^{\beta}}{\prod(x_{i}+y_{j})^{2}}\frac{dx_{1}}{x_{1}}\wedge\cdots\wedge\frac{dy_{\ell}}{y_{\ell}},$
such that in $x^{\alpha}=x^{\alpha_{1}}\cdots x^{\alpha_{k}},$ at least one $\alpha_{i}\leq-2$. The integral with respect to the $x_{j}$ can be broken up into the product of integrals
$\oint\frac{x_{i}^{\alpha_{j}}}{\prod_{m}(x_{j}+y_{m})^{2}}\frac{dx_{j}}{x_{j}}.$
For $j=i$, the fraction will be of degree at most 2-2 and so must be zero by Lemma 8. ∎
Now that we have this inner product, we can reformulate the equations for $m_{\lambda}^{\prime}$, $P(k,\ell;n)$ and $P^{\prime}(k,\ell;n,m)$ in terms of it.
4 Proof of Budzik’s Conjecture
The proof is now easy:
$m_{\lambda}(k,\ell)$ $-m_{\lambda}(k-1,\ell-1)$
$=\sum_{\mu}\{\gamma_{\mu,\mu}|\mu\in H(k,\ell)/H(k-1,\ell-1)\}$
[…]
$=\langle HS_{\lambda}(XX^{-1}\cup YY^{-1};XY^{-1}\cup X^{-1}Y),1\rangle\mbox{ (by Rosas's Theorem)}$
$=\langle HS_{\lambda}(Z_{0};Z_{1}),1\rangle$
$=m_{\lambda}^{\prime}$
5 Concomitants
In [1] we studied not only the invariants of $U(k|\ell)$, but also the concomitants. Each of the three definitions of $m_{\lambda}$ have analogues to multiplicities we call $\bar{m}_{\lambda}$. In terms of tensor products $\sum\bar{m}_{\lambda}$ is the multiplicity of $\chi^{\lambda}$ in $\sum_{\mu\in H(k,\ell)}(\chi^{\mu}\otimes\chi^{\mu})\downarrow,$ where the arrow denotes inducing down from each $S_{n}$ to $S_{n-1}$. For the second one, we use $\bar{F}[X_{1},\ldots,X_{n}]$, defined to be the tensor product $F[X_{1},\ldots,X_{n}]\otimes TR(k,\ell;n)$. This algebra is $n$-graded and has Poincaré series $\sum\bar{m}_{\lambda}s_{\lambda}(t_{1},\ldots,t_{n})$, denoted $\bar{P}(k,\ell;n)$. And, for the third one, we use the algebra of $U(k|\ell)$-invariant functions $M_{k,\ell}^{n}\rightarrow M_{k,\ell}$. This turns out to be isomorphic to the previous, and so also has Poincaré series $\bar{P}(k,\ell;n)$.
There is an analogue of Budzik’s conjecture in this case and it, together with the analogues of equations (3), (4) and (5) turn out to be true. Let $Z=Z_{0}\cup Z_{1}.$ The analogue of $m_{\lambda}^{\prime}$ is
$\bar{m}_{\lambda}^{\prime}=\frac{1}{k!\ell!(2\pi i)^{k+\ell}}\oint_{|y_{i}|=1}\oint_{|x_{i}|=1+\epsilon}HS_{\lambda}(Z_{0};Z_{1})\Delta\sum_{z\in Z}z\frac{dx_{1}}{x_{1}}\wedge\cdots\wedge\frac{dy_{\ell}}{y_{\ell}},$
and the analogue of $P^{\prime}(k,\ell;n)$ is $\bar{P}^{\prime}(k,\ell;n)$ which equals $\frac{1}{k!\ell!(2\pi i)^{k+\ell}}$ times the integral
$\oint_{|y_{i}|=1}\oint_{|x_{i}|=1+\epsilon}\frac{\prod_{i}\prod_{z\in Z_{1}}(1+zt_{i})\sum_{z\in Z}z}{\prod_{i}\prod_{z\in Z_{0}}(1-zt_{i})}\Delta\frac{dx_{1}}{x_{1}}\wedge\cdots\wedge\frac{dy_{\ell}}{y_{\ell}}$
Note that in both cases, the integrals are gotten from those of $m_{\lambda}^{\prime}$ and $P^{\prime}(k,\ell;n)$ by adding in a factor of $\sum_{z\in Z}z$. $P^{\prime}(k,\ell;n,m)$ is defined in the same way, but we don’t write it out because of its length.
To start the proof, we relate $\bar{m}_{\lambda}^{\prime}$ to $\bar{P}^{\prime}(k,\ell;n,m)$ and $\bar{P}^{\prime}(k,\ell;n)$.
Lemma 10.
The functions $\bar{P}^{\prime}(k,\ell;n,m)$ and $\bar{P}^{\prime}(k,\ell;n)$ satisfy $\bar{P}^{\prime}(k,\ell;n)=\sum\bar{m}_{\lambda}^{\prime}s_{\lambda}(T)$ and $\bar{P}^{\prime}(k,\ell;n,m)=\sum m^{\prime}_{\lambda}HS_{\lambda}(T;U)$.
Proof.
For simplicity, we prove only the first statement. The proof of the second is essentially the same.
$\sum_{\lambda}\bar{m}^{\prime}_{\lambda}s_{\lambda}(T)$ $=\sum_{\lambda}\langle HS_{\lambda}(Z_{0};Z_{1})\sum z,1\rangle s_{\lambda}(T)$
$=\langle HS_{\lambda}(Z_{0};Z_{1})s_{\lambda}(T)\sum z,1\rangle$
$=\langle\prod_{i}\prod_{z\in Z_{0}}(1+zt_{i})\prod_{i}\prod_{z\in Z_{1}}(1+zu_{i})\times$
$\quad\prod_{i}\prod_{z\in Z_{1}}(1+zt_{i})^{-1}\prod_{i}\prod_{z\in Z_{0}}(1+zu_{i})^{-1}\sum z\rangle$
$=P^{\prime}(k,\ell;n)$
The third inequality is Theorem 6. ∎
The next lemma is a standard trick in trace identity theory.
Lemma 11.
$\bar{P}(k,\ell;n)$ can be gotten from $P(k,\ell;n+1)$ via
$\bar{P}(k,\ell;n)=\left.\frac{\partial P(k,\ell;n+1)}{\partial t_{n+1}}\right|_{t_{n+1}=0}.$
Proof.
We map $\bar{F}[X_{1},\ldots,X_{n}]\rightarrow TR(k,\ell;n+1)$ via $f\mapsto tr(X_{k+1}f)$. Since the trace is non-degenerate, this is an $n$-grading preserving linear isomorphism onto the part of $TR(k,\ell;n+1)$ of degree 1 in $X_{k+1}$. Hence, the Poincaré series of $\bar{F}[X_{1},\ldots,X_{n}]$ is equal to the Poincaré series of the subspace of $TR(k,\ell;n+1)$ of degree 1 in $X_{k+1}$. The operation of taking derivative with respect to $t_{k+1}$ and then setting $t_{k+1}$ equal to 0 computes the part of $\bar{P}(k,\ell;n+1)$ of degree 1 in $t_{k+1}$. ∎
The next lemma does this computation.
Lemma 12.
For each $i$, $\frac{\partial}{\partial t_{i+1}}P^{\prime}(k,\ell,i+1)|_{t_{i+1}=0}=\bar{P}^{\prime}(k,\ell;i)$.
Proof.
The integrand computing $P^{\prime}$ from Theorem 1, taking $n=i$, is a product of terms not involving $t_{i}$ times $\prod_{z\in Z_{1}}(1+zt_{i})\prod_{z\in Z_{0}}(1-z_{i}t)^{-1}$. Taking the partial derivative with respect to $t_{i}$ and then setting $t_{i}=0$ gives a sum of terms $\sum_{z\in Z_{1}}z+\sum_{z\in Z_{0}}z=\sum_{z\in Z}z.$. ∎
Here is the main theorem of this section.
Theorem 13.
For each $\lambda$, $\bar{m}_{\lambda}^{\prime}(k,\ell)=\bar{m}_{\lambda}(k,\ell)-\bar{m}_{\lambda}(k-1,\ell-1)$, and so $\bar{m}_{\lambda}(k,\ell)=\sum m_{\lambda}^{\prime}(k-i,\ell-i)$.
For each $n$, $\bar{P}^{\prime}(k,\ell;n)=\bar{P}(k,\ell;n)-\bar{P}(k-1,\ell-1;n)$, and so $\bar{P}(k,\ell;n)=\sum\bar{P}^{\prime}(k-i,\ell-i;n)$. And, more generally, $\bar{P}^{\prime}(k,\ell;n,m)=\bar{P}(k,\ell;n,m)-\bar{P}(k-1,\ell-1;n,m)$, and so $\bar{P}(k,\ell;n,m)=\sum\bar{P}^{\prime}(k-i,\ell-i;n,m)$.
Proof.
We begin with the case of $\bar{P}(k,\ell;n).$
[…]
which implies $\bar{P}^{\prime}(k,\ell;n)=\bar{P}(k,\ell;n)-\bar{P}(k-1,\ell-1;n)$.
This implies, using Lemma 10, that
$\sum\bar{m}_{\lambda}^{\prime}(k,\ell)s_{\lambda}=\sum\bar{m}_{\lambda}^{\prime}(k,\ell)s_{\lambda}(k,\ell)-\sum\bar{m}_{\lambda}^{\prime}(k-1,\ell-1)s_{\lambda}(k,\ell)$
and so each $\bar{m}_{\lambda}^{\prime}(k,\ell)=\sum\bar{m}_{\lambda}(k,\ell)-\sum\bar{m}(k-1,\ell-1)$, which in term implies $\bar{P}^{\prime}(k,\ell,n,m)=P(k,\ell;n,m)-P(k-1,\ell-1;n,m)$.
6 Algebraic Properties of the Poincaré Series
Our main goal in this section is to prove that each of the Poincaré series studied in this paper, namely, $P(k,\ell;n)$, $P(k,\ell;n,m)$, $\bar{P}(k,\ell;n)$, $\bar{P}(k,\ell;n,m)$, $P^{\prime}(k,\ell;n)$, $P^{\prime}(k,\ell;n,m)$ $\bar{P}^{\prime}(k,\ell;n)$ and $\bar{P}^{\prime}(k,\ell;n,m)$ is a nice rational function, by which we mean that each is a rational function whose denominator can be written as a product of terms of the form 1w1-w, where $w$ is a monic monomial. We will also prove that the monomials $w$ each have degree at most $k+\ell$. Moreover, we prove that if $\ell=0$, each $w$ has degree at most $\max\{k,\ell\}$; and in any case $w$ has even degree in the $u_{i}$. [^1] The computed value of $\bar{P}^{\prime}(1,1;2,2)$ in [1] has a factor of $(1-u_{1})(1-u_{2})$ in the denominator. However, multiplying numerator and denominator by $(1+u_{1})(1+u_{2})$ gives a fraction with $(1-u_{1}^{2})(1-u_{2}^{2})$ in the denominator, both terms of even degree in $u_{i}$ and both of degree $\leq 2$. Note that the first four in our list (the unprimed ones) are sums of the last four, so it suffices to prove that these four are nice rational. For ease of exposition we will first do the case of $P^{\prime}(k,\ell;n)$ and then indicate the changes needed for the remaining cases.
By Theorem 1, $P^{\prime}(k,\ell;n)$ is the integral of
$\frac{\prod_{i,j,\alpha}(1+x_{i}^{-1}y_{j}t_{\alpha})(1+x_{i}y_{j}^{-1}t_{\alpha})\prod_{i\neq j}(1-x_{i}x_{j}^{-1})\prod_{i\neq j}(1-y_{i}y_{j}^{-1})}{\prod_{i,j,\alpha}(1-x_{i}x_{j}^{-1}t_{\alpha})\prod_{i,j,\alpha}(1-y_{i}y_{j}^{-1}t_{\alpha})\prod_{i}x_{i}\prod_{j}y_{j}}$ (9)
times
$\prod_{i,j}(1+x_{i}^{-1}y_{j})^{-1}(1+x_{i}y_{j}^{-1})^{-1}.$ (10)
We first focus on this latter product. For each $i,j,$ $x_{i}$ has norm $1+\epsilon$ and $y_{j}$ has smaller norm, 1, so we can write each $(1+x_{i}^{-1}y_{j})^{-1}(1+x_{i}y_{j}^{-1})^{-1}$ as $x_{i}^{-1}y_{j}(1+x_{i}^{-1}y_{j})^{-2}=x_{i}^{-1}y_{j}\sum_{m=0}^{\infty}(m+1)(-x_{i}^{-1}y_{j})^{m}.$ Hence, (10) equals $\prod_{i}x_{i}^{-\ell}\prod_{j}y_{j}^{k}\prod_{i,j}\sum_{m=0}^{\infty}(m+1)(-x_{i}^{-1}y_{j})^{m}$.
In (9) the degree in each $x_{i}$ is $n(k+\ell)-1$, so by Lemma 8, the integral of (9) times $(-x^{-1}y_{j})^{m}$ will be zero if $m\geq n(k+\ell)+1$. Letting $N$ equal to $n(k+\ell)+1$, we can replace each $\sum_{m=0}^{\infty}(m+1)(-x_{i}^{-1}y_{j})^{m}$ by $\sum_{m=0}^{N}(m+1)(-x_{i}^{-1}y_{j})^{m}$ without affecting the integral. Hence, $P^{\prime}(k,\ell;n)$ is a linear combination of (many) integrals with integrands of the form
$\frac{x_{1}^{i_{1}}\cdots x_{k}^{i_{k}}y_{1}^{j_{1}}\cdots y_{\ell}^{j_{\ell}}t_{1}^{\alpha_{1}}\cdots t_{n}^{\alpha_{n}}}{\prod_{i,j,\alpha}(1-x_{i}x_{j}^{-1}t_{\alpha})\prod_{i,j,\alpha}(1-y_{i}y_{j}^{-1}t_{\alpha})}.$ (11)
That such an integral is a nice rational function follows directly from Theorem 3.4 of [9], which in turn is based on Stanley’s Corollary 3.8 in [8]. Van Den Bergh’s theory associates directed graphs to integrals of this form. In our case, the graph would have vertices $v[x_{i}]$ and $v[y_{j}]$, corresponding to the variables $x_{1},\ldots,y_{\ell}$, and with an edge $e[x_{i},x_{j},t_{\alpha}]$ connecting $v[x_{i}]$ with $v[x_{j}]$ for each term $(1-x_{i}x_{j}^{-1}t_{\alpha})$ in the denominator, and an edge $e[y_{i},y_{j},t_{\alpha}]$ from $v[y_{i}]$ to $v[y_{j}]$ for each $(1-y_{i}y_{j}^{-1}t_{\alpha})$. Then his Theorem 3.4 not only states that the integral is a nice rational function, but identitifies the denominator has a product of terms of the form $1-t^{C}$, where $C$ is an oriented cycle in the graph, and where $t^{C}$ means the product of the $t_{\alpha}$ corresponding to edges in $C$. Since a cycle cannot have repeated vertices, and since edges connect only the $v[x_{i}]$ or the $v[y_{j}]$, but not both, $|C|\leq\max\{k,\ell\}$, and so the degree of $t^{C}$ is at most $\max\{k,\ell\}$.
The case of $\bar{P}^{\prime}(k,\ell;n)$ is essentially the same. As for the cases of $P^{\prime}(k,\ell;n,m)$ and $\bar{P}^{\prime}(k,\ell;n,m)$, the integrands (11) would have extra factors of
$\prod(1-x_{i}y_{j}^{-1}u_{\alpha})(1-x_{i}^{-1}y_{j}u_{\alpha}).$
In terms of the graph theory, we would add to the previous graph edges $e[x_{i},y_{j},u_{\alpha}]$ for each $(1-x_{i}y_{j}^{-1}u_{\alpha})$ and $e[y_{i},x_{j},u_{\alpha}]$ for each $(1-x_{i}^{-1}y_{j}u_{\alpha})$. The effect is that cycles could now have length as great as $k+\ell$, the number of vertices, and the degree in the edges corresponding to the $u_{\alpha}$ must be even.

### v2

@@
 Abstract
-In [1] we defined integrals that approximated the Poincaré series of the invariants and cocomiants of the general linear Lie supergroup or superalgebra.
+In [2] we defined integrals that approximated the Poincaré series of the invariants and concomiants of the general linear Lie supergroup or superalgebra.
 Budzik suggested in [4] a way to adapt this method to get the exact Poincaré series.
@@
 1 Budzik’s Conjecture
-The basic objects of that form the subject of this paper can be described in several different ways.
+The basic objects that form the subject of this paper can be described in several different ways.
 The proof of our main theorem will mosty focus on their definition based on the character theory of the symmetic group, but in order to provide motivation and context, we also mention some of the connections to trace identities and to invariant theory.
@@
 $m_{\lambda}=\sum_{\mu\in H(k,\ell)}\gamma_{\mu,\mu}^{\lambda}.$ (1)
-Sometimes we will write $m_{\lambda}(k,\ell)$ if we need to specify the $k$ and $\ell$.
-We associate to $m_{\lambda}$ the Poincaré series $P(k,\ell;n)=\sum m_{\lambda}s_{\lambda}(t_{1},\ldots,t_{n})$, where $s_{\lambda}$ is the Schur function, and more generally,
+Sometimes we will write $m_{\lambda}(k,\ell)$ if we need to specify the $k$ and $\ell$. We associate to $m_{\lambda}$ the Poincaré series
+$P(k,\ell;n)=\sum m_{\lambda}s_{\lambda}(t_{1},\ldots,t_{n}),$
+where $s_{\lambda}$ is the Schur function, and more generally,
 $P(k,\ell;n,m)=\sum m_{\lambda}HS_{\lambda}(t_{1},\ldots,t_{n};u_{1},\ldots,u_{m}),$
@@
 for all $a_{1},\ldots,a_{n}\in M_{k,\ell}$ and all $g\in U(k|\ell)$.
-The algebra of invariant functions on $M_{k,\ell}^{n},$ polynomial in the entries, also has an $n$-fold grading and also has Poincaré series $P(k,\ell,n)$.
+The algebra of invariant functions on $M_{k,\ell}^{n},$ polynomial in the entries, also has an $n$-fold grading and also has Poincaré series $P(k,\ell,n)$, the same as above.
 In fact it is generated by the trace polynomials.
-The series $P(k,\ell;n,m)$ also has a natural interpretation.
-We refer the interested reader to [1].
-We also mention that we learned from [3] that these series have signifigance in physics: “These arise in physics e.g.
-as free partition functions of $U(N|M)$ gauge theories or supersymmetric indices.”
+The series $P(k,\ell;n,m)$ also has a natural interpretation, see [1].
+The algebra of $(k+\ell)\times(k+\ell)$ matrices with entries in $E$ has a $\mathbb{Z}_{2}$-grading in which $M_{k,\ell}$ is the degree 0 part.
+The degree 1 part would consist of matrices $\left(\begin{smallmatrix}A&B\\ C&D\end{smallmatrix}\right)$,with dimensions as above, but in which $A$ and $D$ have degree 1 entries and $B$ and $C$ have degree 0 ones.
+This algebra has a supertrace function, defined on a homogeneous matrix as above as $tr(A)\mp tr(D)$, depending on whether the matrix is degree 0 or 1.
+Letting $F[X_{1},\ldots,X_{n},Y_{1},\ldots,Y_{m}]$ be the algebra generated by $n$ generic degree 0 matrices and $m$ generic degree 1 matrices, we can then let $STR(k,\ell;n,m)$ be the algebra generated by the supertraces.
+This algebra has an $(n+m)$-fold grading and its Poincaré series is $P(k,\ell;n,m)$.
+We also mention that we learned from [4] that these series have signifigance in physics: “These arise in physics e.g. as free partition functions of $U(N|M)$ gauge theories or supersymmetric indices.”
 In [2] we studied these Poincaré series using the complex integrals. Let $\Delta$ (not the same as the $\Delta$ from [2]) be defined as
@@
 $m_{\lambda}(k,\ell)=\sum_{i=0}^{\min(k,\ell)}m_{\lambda}^{\prime}(k-i,\ell-i).$ (3)
-Likewise, setting $P^{\prime}(k,\ell;n)$ equal to $\sum m_{\lambda}^{\prime}S_{\lambda}(t_{1},\ldots,t_{n})$, Budzik’s conjecture would imply that $P^{\prime}(k,\ell;n)=P(k,\ell;n)-P(k-1,\ell-1;n)$ and so
+Likewise, since $P^{\prime}(k,\ell;n)$ is equal to $\sum m_{\lambda}^{\prime}S_{\lambda}(t_{1},\ldots,t_{n})$, Budzik’s conjecture would imply that $P^{\prime}(k,\ell;n)=P(k,\ell;n)-P(k-1,\ell-1;n)$ and so
 $P(k,\ell;n)=\sum_{i=0}^{\min(k,\ell)}P^{\prime}(k-i,\ell-i;n)$ (4)
-More generally, setting $P^{\prime}(k,\ell;n,m)$ equal to $\sum m_{\lambda}^{\prime}HS(t_{1},\ldots,t_{n};u_{1},\ldots,u_{m})$ we would have $P^{\prime}(k,\ell;n,m)=P(k,\ell;n,m)-P(k-1,\ell-1k;n)$ and
+More generally, since $P^{\prime}(k,\ell;n,m)$ is equal to $\sum m_{\lambda}^{\prime}HS_{\lambda}(t_{1},\ldots,t_{n};u_{1},\ldots,u_{m})$ we would have $P^{\prime}(k,\ell;n,m)=P(k,\ell;n,m)-P(k-1,\ell-1;n,m)$ and
 $P(k,\ell;n,m)=\sum_{i=0}^{\min(k,\ell)}P^{\prime}(k-i,\ell-i;n,m).$ (5)
@@
 and, taking $Y=\emptyset$, $P^{\prime}(k,\ell;n)$ equals $\frac{1}{k!\ell!(2\pi i)^{k+\ell}}$ times the integral
-$\oint_{|y_{i}|=1}\oint_{|x_{i}|=1+\epsilon}\frac{\prod_{i}\prod_{z\in Z_{1}}(1+zt_{i})}{\prod_{i}\prod_{z\in Z_{0}}(1-zt_{i})\prod_{i}}\Delta\frac{dx_{1}}{x_{1}}\wedge\cdots\wedge\frac{dy_{\ell}}{y_{\ell}}$
-Before leaving this section we also mention that Budzik computed $P^{\prime}(k,\ell;1)$ for all $k$ and $\ell$, namely,
-$P^{\prime}(k,\ell;1)=\frac{t^{k\ell}}{\prod_{i=1}^{k}(1-t^{i})\prod_{j=1}^{\ell}(1-t^{j})}.$
+$\oint_{|y_{i}|=1}\oint_{|x_{i}|=1+\epsilon}\frac{\prod_{i}\prod_{z\in Z_{1}}(1+zt_{i})}{\prod_{i}\prod_{z\in Z_{0}}(1-zt_{i})}\Delta\frac{dx_{1}}{x_{1}}\wedge\cdots\wedge\frac{dy_{\ell}}{y_{\ell}}$
 2 Properties of Hook Schur Functions
@@
 For each partition $\lambda$ we let $HS_{\lambda}(X;Y)$ be the hook Schur function on $\lambda$.
-This is commonly defined as $\sum_{\mu\subseteq\lambda}s_{\lambda}(X)s_{\lambda/\mu}(Y)$, where $s$ is the ordinary Schur function.
+This is commonly defined as $\sum_{\mu\subseteq\lambda}s_{\mu}(X)s_{\lambda/\mu}(Y)$, where $s$ is the ordinary Schur function.
 Some authors study the super Schur funcitons $s_{\lambda}(X/Y)$ instead of the hook Schur functions $HS_{\lambda}(X;Y)$.
@@
 Then $f_{\lambda}$ is defined to be the product $\prod(x_{i}+y_{j})$ over all $(i,j)\in\lambda$, with the convention that $x_{i}=0$ and $y_{j}=0$ for $i>k$ or $j>\ell$.
-For example, if $(k,\ell)=(2,1)$ and $\lambda=(2,1)$, then $f_{\lambda}=(x_{1}+y_{1})x_{1}^{2}(x_{2}+y)$.
+For example, if $(k,\ell)=(2,1)$ and $\lambda=(2,1)$, then $f_{\lambda}=(x_{1}+y_{1})x_{1}(x_{2}+y)$.
 Józefiak and Pragacz’s theorem gives an alternate description of hook Schur functions.
@@
 Theorem 4 (Berele and Regev).
-Let $\lambda$ be a typical partition, and define the partitions $\alpha$ and $\beta$, via $\alpha=(\lambda_{1}-\ell,\ldots,\lambda_{k}-\ell)$ and $\beta=(\lambda_{1}^{\prime}-k,\ldots,\lambda_{\ell}^{\prime}-k)$.
-Then $HS_{\lambda}(X;Y)=\prod_{i,j}(x_{i}+y_{j})s_{\alpha}(X)s_{\beta}(Y)$.
-The next theorem is due to Rosas in [7]. It gives a hook Schur function interpretation of the coefficients $\gamma_{\mu\nu}^{\lambda}$ occuring in the tensor product of the symmetric group characters.
+Let $\lambda$ be a typical partition, and define the partitions $\mu$ and $\nu$, via $\mu=(\lambda_{1}-\ell,\ldots,\lambda_{k}-\ell)$ and $\nu=(\lambda_{1}^{\prime}-k,\ldots,\lambda_{\ell}^{\prime}-k)$, see Figure 1.
+Then $HS_{\lambda}(X;Y)=\prod_{i,j}(x_{i}+y_{j})s_{\mu}(X)s_{\nu}(Y)$.
+The next theorem is due to Rosas in [10].
+It gives a hook Schur function interpretation of the coefficients $\gamma_{\mu\nu}^{\lambda}$ occuring in the tensor product of the symmetric group characters.
 Theorem 5 (Rosas).
@@
 Lemma 7.
-If $\mu$ and $\nu$ are not both in $H^{\prime}(k,\ell)$, then $\langle HS_{\mu},HS_{\nu}\rangle=0$.
+If $\mu$ and $\nu$ are not both typical, then $\langle HS_{\mu},HS_{\nu}\rangle=0$.
 If $\mu$ or $\nu$ are not in $H(k,\ell)$, then $H_{\mu}(X;Y)$ or $H_{\nu}(X;Y)$ will be zero, so we will focus on $\mu,\nu\in H(k,\ell)$.
@@
 Let $f=f(z_{1},\ldots,z_{n})$ and $g=g(z_{1},\ldots,z_{n})$ be Laurant polynomials with leading terms of degrees $a$ and $b$, respectively, with a<b1a<b-1.
-Let $|z|=r$ be a circle containing all of the roots of $g(z)$.
+Let $|z|=r$ be a circle containing all of the roots of $g(z)$ in its interior.
 Then $\oint_{|z|=r}f(z)/g(z)dz=0$.
@@
 $\sigma\left[f_{\mu}(X,Y)\prod_{i<j}(1-x_{j}x_{i}^{-1})^{-1}\prod_{i<j}(1-y_{j}y_{i}^{-1})^{-1}\right].$ (8)
-Since $\mu$ is not typical, $\mu_{k}\leq\ell-1$, so $f_{\mu}(X;Y)$ is of degree at most $\ell-1$ in $x_{k}$; also $\prod_{i<j}(1-x_{j}x_{i}^{-1})^{-1}$ will be of degree k1k-1 in $x_{k}$, so (8) will be of degree at most $\ell-k$ in $x_{k}$, and (6) will be of degree at most 2.
+Since $\mu$ is not typical, $\mu_{k}\leq\ell-1$, so $f_{\mu}(X;Y)$ is of degree at most $\ell-1$ in $x_{k}$; also $\prod_{i<j}(1-x_{j}x_{i}^{-1})^{-1}$ will be of degree k1k-1 in $x_{k}$, so (8) will be of degree at most $\ell-k$ in $x_{k}$.
 Under a permutation $\sigma$, $f_{\lambda}(X;Y)/\prod(1-x_{j}x_{i}^{-1})$ will be of degree at most $\ell-k$ in $x_{\sigma(k)}$, and the product with (7) at most 2-2.
@@
 $\oint_{|y|=1}\oint_{|x|=1+\epsilon}\frac{x^{\alpha}y^{\beta}}{\prod(x_{i}+y_{j})^{2}}\frac{dx_{1}}{x_{1}}\wedge\cdots\wedge\frac{dy_{\ell}}{y_{\ell}},$
-such that in $x^{\alpha}=x^{\alpha_{1}}\cdots x^{\alpha_{k}},$ at least one $\alpha_{i}\leq-2$. The integral with respect to the $x_{j}$ can be broken up into the product of integrals
-$\oint\frac{x_{i}^{\alpha_{j}}}{\prod_{m}(x_{j}+y_{m})^{2}}\frac{dx_{j}}{x_{j}}.$
+such that for at least one $i$, $x_{i}^{\alpha_{i}}\prod_{j}(x_{i}+y_{j})^{-2}$the degree in $x_{i}$ is less than or equal to 2-2.
+The integral with respect to the $x_{j}$ can be broken up into the product of integrals
+$\oint\frac{x_{j}^{\alpha_{j}}}{\prod_{m}(x_{j}+y_{m})^{2}}\frac{dx_{j}}{x_{j}}.$
 For $j=i$, the fraction will be of degree at most 2-2 and so must be zero by Lemma 8. ∎
-Now that we have this inner product, we can reformulate the equations for $m_{\lambda}^{\prime}$, $P(k,\ell;n)$ and $P^{\prime}(k,\ell;n,m)$ in terms of it.
 4 Proof of Budzik’s Conjecture
@@
 […]
-[…]
+$=\langle\sum_{\mu,\nu}\gamma_{\mu,\nu}^{\lambda}HS_{\mu}(X;Y)HS_{\nu}(X^{-1};Y^{-1}),1\rangle$
 $=\langle HS_{\lambda}(XX^{-1}\cup YY^{-1};XY^{-1}\cup X^{-1}Y),1\rangle\mbox{ (by Rosas's Theorem)}$
@@
 $=m_{\lambda}^{\prime}$
+Note that using Theorem 6 we can directly compute the formula for $P^{\prime}(k,\ell;n,m)$:
+$P^{\prime}(k,\ell;n,m)$ $=\sum m_{\lambda}HS_{\lambda}(T,U)$
+$=\sum\langle HS_{\lambda}(Z_{0};Z_{1}),1\rangle HS_{\lambda}(T;U)$
+$=\langle\sum HS_{\lambda}(Z_{0};Z_{1})HS_{\lambda}(T;U),1\rangle$
+The sum is the integrand of Theorem  by the Berele-Remmel Theorem, Theorem 6.
 5 Concomitants
@@
 This turns out to be isomorphic to the previous, and so also has Poincaré series $\bar{P}(k,\ell;n)$.
+Likewise for $\bar{P}(k,\ell;n,m)$.
 There is an analogue of Budzik’s conjecture in this case and it, together with the analogues of equations (3), (4) and (5) turn out to be true.
@@
 $=\langle\prod_{i}\prod_{z\in Z_{0}}(1+zt_{i})\prod_{i}\prod_{z\in Z_{1}}(1+zu_{i})\times$
-$\quad\prod_{i}\prod_{z\in Z_{1}}(1+zt_{i})^{-1}\prod_{i}\prod_{z\in Z_{0}}(1+zu_{i})^{-1}\sum z\rangle$
+$\quad\prod_{i}\prod_{z\in Z_{1}}(1+zt_{i})^{-1}\prod_{i}\prod_{z\in Z_{0}}(1+zu_{i})^{-1}\sum z,1\rangle$
 $=P^{\prime}(k,\ell;n)$
@@
 Lemma 12.
-For each $i$, $\frac{\partial}{\partial t_{i+1}}P^{\prime}(k,\ell,i+1)|_{t_{i+1}=0}=\bar{P}^{\prime}(k,\ell;i)$.
+For each $i$, $\frac{\partial}{\partial t_{i+1}}P^{\prime}(k,\ell;i+1)|_{t_{i+1}=0}=\bar{P}^{\prime}(k,\ell;i)$.
 Proof.
@@
 This implies, using Lemma 10, that
-$\sum\bar{m}_{\lambda}^{\prime}(k,\ell)s_{\lambda}=\sum\bar{m}_{\lambda}^{\prime}(k,\ell)s_{\lambda}(k,\ell)-\sum\bar{m}_{\lambda}^{\prime}(k-1,\ell-1)s_{\lambda}(k,\ell)$
-and so each $\bar{m}_{\lambda}^{\prime}(k,\ell)=\sum\bar{m}_{\lambda}(k,\ell)-\sum\bar{m}(k-1,\ell-1)$, which in term implies $\bar{P}^{\prime}(k,\ell,n,m)=P(k,\ell;n,m)-P(k-1,\ell-1;n,m)$.
+$\sum\bar{m}_{\lambda}^{\prime}(k,\ell)s_{\lambda}=\sum\bar{m}_{\lambda}^{\prime}(k,\ell)s_{\lambda}-\sum\bar{m}_{\lambda}^{\prime}(k-1,\ell-1)s_{\lambda}.$
+Since the Schur functions are linearly independent, this implies that $\bar{m}_{\lambda}^{\prime}(k,\ell)=\sum\bar{m}_{\lambda}(k,\ell)-\sum\bar{m}(k-1,\ell-1)$, which in term implies $\bar{P}^{\prime}(k,\ell,n,m)=P(k,\ell;n,m)-P(k-1,\ell-1;n,m)$.
 6 Algebraic Properties of the Poincaré Series
@@
 For ease of exposition we will first do the case of $P^{\prime}(k,\ell;n)$ and then indicate the changes needed for the remaining cases.
-By Theorem 1, $P^{\prime}(k,\ell;n)$ is the integral of
+By Theorem 4, $P^{\prime}(k,\ell;n)$ is the integral of the product of
 $\frac{\prod_{i,j,\alpha}(1+x_{i}^{-1}y_{j}t_{\alpha})(1+x_{i}y_{j}^{-1}t_{\alpha})\prod_{i\neq j}(1-x_{i}x_{j}^{-1})\prod_{i\neq j}(1-y_{i}y_{j}^{-1})}{\prod_{i,j,\alpha}(1-x_{i}x_{j}^{-1}t_{\alpha})\prod_{i,j,\alpha}(1-y_{i}y_{j}^{-1}t_{\alpha})\prod_{i}x_{i}\prod_{j}y_{j}}$ (9)
-times
+and
 $\prod_{i,j}(1+x_{i}^{-1}y_{j})^{-1}(1+x_{i}y_{j}^{-1})^{-1}.$ (10)
@@
 In (9) the degree in each $x_{i}$ is $n(k+\ell)-1$, so by Lemma 8, the integral of (9) times $(-x^{-1}y_{j})^{m}$ will be zero if $m\geq n(k+\ell)+1$.
-Letting $N$ equal to $n(k+\ell)+1$, we can replace each $\sum_{m=0}^{\infty}(m+1)(-x_{i}^{-1}y_{j})^{m}$ by $\sum_{m=0}^{N}(m+1)(-x_{i}^{-1}y_{j})^{m}$ without affecting the integral.
+Letting $N$ equal to $n(k+\ell)+1$, we can replace each $\sum_{m=0}^{\infty}(m+1)(-x_{i}^{-1}y_{j})^{m}$ by $\sum_{m=0}^{N}(m+1)(-x_{i}^{-1}y_{j})^{m}$ without effecting the integral.
 Hence, $P^{\prime}(k,\ell;n)$ is a linear combination of (many) integrals with integrands of the form
@@
 Then his Theorem 3.4 not only states that the integral is a nice rational function, but identitifies the denominator has a product of terms of the form $1-t^{C}$, where $C$ is an oriented cycle in the graph, and where $t^{C}$ means the product of the $t_{\alpha}$ corresponding to edges in $C$.
-Since a cycle cannot have repeated vertices, and since edges connect only the $v[x_{i}]$ or the $v[y_{j}]$, but not both, $|C|\leq\max\{k,\ell\}$, and so the degree of $t^{C}$ is at most $\max\{k,\ell\}$.
+Since a cycle cannot have repeated vertices, and since edges connect only the $v[x_{i}]$ to each other or the $v[y_{j}]$ to each other, $|C|\leq\max\{k,\ell\}$, and so the degree of $t^{C}$ is at most $\max\{k,\ell\}$.
 The case of $\bar{P}^{\prime}(k,\ell;n)$ is essentially the same. As for the cases of $P^{\prime}(k,\ell;n,m)$ and $\bar{P}^{\prime}(k,\ell;n,m)$, the integrands (11) would have extra factors of
@@
 In terms of the graph theory, we would add to the previous graph edges $e[x_{i},y_{j},u_{\alpha}]$ for each $(1-x_{i}y_{j}^{-1}u_{\alpha})$ and $e[y_{i},x_{j},u_{\alpha}]$ for each $(1-x_{i}^{-1}y_{j}u_{\alpha})$.
-The effect is that cycles could now have length as great as $k+\ell$, the number of vertices, and the degree in the edges corresponding to the $u_{\alpha}$ must be even.
+The effect of these additional edges would be that cycles could now have length as great as $k+\ell$, the number of vertices, and the degree in the edges corresponding to the $u_{\alpha}$ must be even.
+7 Computation of $P^{\prime}(k,\ell;0,1)$
+In [4] Budik computed $P^{\prime}(k,\ell,1,0)$, the Poincaré series of the ring generated by the traces of a single generic element of $M_{k,\ell}$:
+$P^{\prime}(k,\ell;1)=\frac{t^{k\ell}}{\prod_{i=1}^{k}(1-t^{i})\prod_{j=1}^{\ell}(1-t^{j})},$ (12)
+Also computed by Molev and Mukhin in [7], Orellana and Zabrocki in [8].
+A different formula for $P(k,\ell;n)$ was given by Sergeev and Veselov in [11].
+We give yet another derivation based on Theorem 5.2 of [2], our proof being completely combinatorial.
+Theorem 14 (=Theorem 5.2 of [2]).
+$m_{(n)}(k,\ell)$ equals the number of partitions of $n$ in the hook $H(k,\ell)$; and $m_{(1^{n})}(k,\ell)$ equals the number of self-conjugate partitions of $n$ in $H(k,\ell)$.
+It follows that $m^{\prime}_{(n)}(k,\ell)=m_{(n)}(k,\ell)-m_{(n)}(k-1,\ell-1)$ equals the number of partitions in $H(k,\ell)$, but not in $H(k-1,\ell-1)$.
+These are the typical partitions, $H^{\prime}(k,\ell)$.
+A Young diagram of a typical partition consists of three parts: The $k\times\ell$ rectangle, a partition of height at most $k$ to the right of it, and the conjugate of a partition of height at most $\ell$ below it.
+See figure 1.
+Since the generating function for partitions of height at most $k$ is $\prod_{i=1}^{k}(1-x^{i})^{-1}$, (12) follows.
+$k\times\ell$$\mu$$\nu^{\prime}$
+Figure 1: A Typical Partition
+We now turn our attention to $P^{\prime}(k,\ell;0,1)$.
+We will take $k\geq\ell$.
+This is the Poincaré series of the supertrace ring generated by one matrix of the form $\left(\begin{smallmatrix}A&B\\ C&D\end{smallmatrix}\right)$, where $A$ is a $k\times k$ matrix, $D$ is an $\ell\times\ell$ matrix, $A$ and $D$ have generic odd Grassmann entries, and $B$ and $C$ have entries commuting indeterminants.
+As above,
+$m_{(1^{n})}^{\prime}(k,\ell)=m_{(1^{n})}(k,\ell)-m_{(1^{n})}(k-1,\ell-1),$
+which, by Theorem 14, equals the number of self-conjugate partitions of $n$ in the hook $H(k,\ell)$, but not in the hook $H(k-1,\ell-1)$.
+If a self-conjugate partition $\lambda$ lies in the hook $H(k,\ell)$, it will also lie in $H(\ell,k)$, meaning that $\lambda_{k+1}\leq\ell$ and $\lambda_{\ell+1}\leq k$.
+In terms of Young diagrams, such a $\lambda$ can be broken into three parts.
+Definition.
+Given a partition $\lambda$ we define $\lambda_{0}$ to be partition $\lambda\cap(k^{k})$, namely, the partition whose Young diagram is the intersection of the Young diagram of $\lambda$ with that of the square $(k^{k})$; $\mu=\mu(\lambda)$ to be the partition to the right of $\lambda_{0}$; and $\nu=\nu(\lambda)$ to be the conjugate of the partition below it, see Figure 2.
+Note that $\lambda$ is self-conjugate if and only if $\lambda_{0}$ is self-conjugate and $\mu(\lambda)=\nu(\lambda)$.
+Lemma 15.
+The correspondence $\lambda\rightarrow(\lambda_{0},\mu(\lambda))$ is a one-to-one correspondence between self-conjugate partitions in $H^{\prime}(k,\ell)$ and pairs of partitions, where $\mu$ is of height at most $\ell$ and $\lambda_{0}$ is a self-conjugate partition containing $(k^{\ell},\ell^{k-\ell})$ and contained in $(k^{k})$.
+Proof.
+If $\lambda\notin H(k-1,\ell-1)$ then $\lambda_{k}\geq\ell$ and, since $\lambda$ is self-conjugate, $\lambda_{\ell}\geq k$.
+The latter implies that the first $\ell$ rows of $\lambda_{0}$ must equal $k$, and the former that the first $\ell$ columns also equal $k$.
+Hence, $\lambda_{0}$ contains $(k^{\ell},\ell^{k-\ell})$.
+The condition $\lambda\in H(k,\ell)\cap H(\ell,k)$ implies that $\mu$ has height at most $\ell$, and the conditions $\lambda_{\ell}\geq k$ and $\lambda_{k}\geq\ell$ imply that any partition of height at most $\ell$ is possible.
+∎
+$\lambda_{0}$$\nu$$\mu\ell\ellkkkk$
+Figure 2: Construction of $\lambda_{0}$, $\mu$ and $\nu$
+Theorem 16.
+$P^{\prime}(k,\ell;0,1)=u^{2k\ell-\ell^{2}}\prod_{i=1}^{k-\ell}(1+u^{2i-1})\prod_{i=1}^{\ell}(1-u^{2i})^{-1}$.
+Proof.
+The generating function for partitions of height at most $\ell$ is $\prod_{i=1}^{\ell}(1-u^{i})^{-1}$, so the generating function for equal pairs of them is $\prod_{i=1}^{\ell}(1-u^{2i})^{-1}$.
+As for $\lambda_{0}$, $\lambda_{0}$ is made up of two parts: $(k^{\ell},\ell^{k-\ell})$ and a self-conjugate partition contained in the lower right $(k-\ell)\times(k-\ell)$ square.
+The former has degree $k\ell+\ell(k-\ell)=2k\ell-\ell^{2}$, and the latter has generating function $\prod_{i=0}^{k-\ell}(1+u^{2i-1}).$
+Budzik related the formula for $P^{\prime}(k,\ell;n)$ to various important combinatorial identities.
+The same came be done for $P^{\prime}(k,\ell;0,1)$.
+The simplest is based on $k=\ell$, in which case
+$P^{\prime}(k,k;0,1)=\frac{u^{k^{2}}}{[u^{2}]_{k}},$
+the generating function for self-conjugate partitions in $H^{\prime}(k,k)$. Using $P(k,k;0,1)=\sum_{n=0}^{k}P^{\prime}(n,n;0,1)$ and letting $k$ go to infinity we get the well-known identity
+$\prod_{n=1}^{\infty}(1+u^{2n-1})=\sum_{k=0}^{\infty}\frac{u^{k^{2}}}{[u^{2}]_{k}}$
+since $P(\infty,\infty;0,1)$ is the generating function for all self-conjugate partitions.
+Similarly, using the identity $P(n+\ell,\ell;0,1)=\sum_{i=0}^{\ell}P^{\prime}(n+i,i;0,1)$ and letting $\ell$ go to infinity, we get
+$\prod_{i=1}^{\infty}(1+u^{2i-1})=\sum_{i=0}^{\infty}\frac{u^{i(2n+i)}\prod_{t=1}^{i}(1+u^{2t-1})}{[u^{2}]_{i}}.$